OpenStudy (anonymous):

Find the x – intercept of -4x + y = -3

OpenStudy (bahrom7893):

same as before, come on!

OpenStudy (bahrom7893):

set y=0, solve for x!

OpenStudy (anonymous):

sorry :( i'm not smart in this field..SOOOO...How do i solve for x?

OpenStudy (bahrom7893):

-4x+0= -3

OpenStudy (anonymous):

-4x + 0 = -3 -4x = -3 divide by sides by -4 x = 3/4

OpenStudy (anonymous):

-3/4?

OpenStudy (anonymous):

not negative because it was -3/-4 so the answer is 3/4

OpenStudy (anonymous):